4. There are 3 green cards, 4 red cards and 5 brown cards in a bag. What is the probability
a.) of getting a red card or a brown card after a green one (w/o replacement)?

b.) that a brown card is drawn first but not replaced, and a red card follows at the second draw?

Answer by math_tutor2020(3835)


Part (a)

3 green + 4 red + 5 brown = 12 total
3/12 = 1/4 represents the probability of getting a green card.

Since the card is not put back, there are 12-1 = 11 cards left.

4 red + 5 brown = 9 cards we want out of 11 left over.
9/11 represents the probability of getting either red or brown on the second selection

(1/4)*(9/11) = 9/44 is the probability of getting green first, then red or brown second, where no replacement is made.

Answer: 9/44

============================================================
Part (b)

5 brown out of 12 total
5/12 = probability of getting a brown card

12-1 = 11 cards left
There are 4 red out of 11 left over
4/11 = probability of getting a red card after the first card is not replaced

(5/12)*(4/11) = (5*4)/(12*11)
(5/12)*(4/11) = (5*4)/(4*3*11)
(5/12)*(4/11) = 5/(3*11)
(5/12)*(4/11) = 5/33
This is the probability of getting brown first, then red next, assuming no replacements are made.

Answer: 5/33